The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Eva K. Grebel and You-Hua Chu

                                    mcnews@astro.uiuc.edu
                    http://www.astro.uiuc.edu/projects/mcnews/MCNews.html
                            http://www.astro.uni-bonn.de/~mcnews/

No. 46 August 6, 2000

Contents

News  1
Abstracts of 10 refereed papers 2
Abstract of a conference paper 9
Job advertisement  9

News

MCNews’ publication policy on the web

We are pleased to announce that a draft publication policy and FAQ for the Magellanic Clouds Newsletter are now available at:

               http://www.astro.uiuc.edu/projects/mcnews/MCpolicy.html

and at the European mirror site at:

                 http://www.astro.uni-bonn.de/~mcnews/MCpolicy.html

These webpages summarize the rules that we adopt in running this newsletter and specify what we publish, how we publish, how to subscribe, etc. Many of these items have appeared earlier in MCNews editorials. The “policy” pages are an attempt to gather everything in one place to make it easier for interested parties to check out our current practice in case questions arise. We hope that you will find this compilation useful.

Eva Grebel & You-Hua Chu

Abstracts of Refereed Papers

Metal Abundances in the Magellanic Stream

Brad K. Gibson^{1}, Mark L. Giroux^{1}, Steven V. Penton^{1}, Mary E. Putman^{2}, John T. Stocke^{1} and J. Michael Shull^{1}

1 Center for Astrophysics & Space Astronomy, University of Colorado, Boulder, CO, 80309-0389, USA 2 RSAA, Australian National University, Weston Creek P.O., Weston, ACT, 2611, Australia

We report on the first metallicity determination for gas in the Magellanic Stream, using archival HST GHRS data for the background targets Fairall 9, III Zw 2, and NGC 7469. For Fairall 9, using two subsequent HST revisits and new Parkes Multibeam Narrowband observations, we have unequivocally detected the MS I H i component of the Stream (near its head) in S ii λλ1250,1253 yielding a metallicity of [S ii/H]= −0.55±0.06(r)_{−0.21}^{+0.17}(s), consistent with either an SMC or LMC origin and with the earlier upper limit set by Lu, Savage & Sembach (1994). We also detect the saturated Si ii λ1260 line, but set only a lower limit of [Si ii/H]∼ >−1.5. We present two HST serendipitous detections of the Stream, seen in Mg ii λλ2796,2803 absorption with column densities of (0.5−1)×10^{13} cm^{−2} toward the Seyfert galaxies III Zw 2 and NGC 7469. These latter sightlines probe gas near the tip of the Stream (∼15^{∘} from the peak of the MS V H i component and ∼80^{∘} down-Stream of Fairall 9). In the case of III Zw 2, the lack of an accurate H i column density determination and the uncertain Mg iii ionization correction severely limit the degree to which we can constrain [Mg/H]; we found a lower limit of [Mg ii/H i]∼ >−1.3 for this sightline. For NGC 7469, an accurate H i column density determination exists, but the extant FOS spectrum limits our ability to constrain the Mg ii column density, and we conclude that [Mg ii/H i]∼ >−1.5 for this sightline. Ionization corrections associated with Mg iii and H ii suggest that the corresponding [Mg/H] may range lower by ∼0.3−1.0 dex. However, an upward revision of ∼0.5−1.0 dex would be expected under the assumption that the Stream exhibits a dust depletion pattern similar to that seen in both the Large and Small Magellanic Clouds. While our abundance analysis allows us to rule out a primordial origin for the Stream, the remaining systematic uncertainties in the H i column density along the lines of sight makes it difficult to differentiate between an LMC versus SMC origin.

Accepted by: The Astronomical Journal

For preprints, contact bgibson@casa.colorado.edu
Also available from the URL http://casa.colorado.edu/∼bgibson/publications.html

Detection of mid-infrared Aromatic Hydrocarbon Emission Features from the Small Magellanic Cloud

William T. Reach^{1,2}, François Boulanger^{2}, Alessandra Contursi^{1}, James Lequeux^{3}

1 Infrared Processing and Analysis Center, California Institute of Technology, MS 100-22, Pasadena, CA 91125, USA 2 Institut d’Astrophysique Spatiale, Bât. 121, Université Paris XI, F–91405 Orsay cedex, France 3 Observatoire de Paris, 61 Av. de l’Observatoire, F-75014 Paris, France

The mid-infrared (5–16 μm) spectral energy distribution for an individual quiescent molecular cloud in the Small Magellanic Cloud (SMC) was observed using ISOCAM. The spectrum is dominated by broad emission bands at 6.2, 7.7, and 11.3 μm, with weaker bands at 8.6 and 12.7 μm. As these are the same bands, with similar shape and relative strengths, as observed in the ISM of our Galaxy, the same carriers must exist in both galaxies. The carriers are widely considered to be large molecules or clusters of aromatic hydrocarbons, which absorb ultraviolet and visible photons and emit mid-infrared photons during high-temperature pulses. Based on the brightness of the mid-infrared emission and the estimated strength of the radiation field in the SMC, the absorption by aromatic hydrocarbons is of order 10% of total dust absorption, comparable to the case for Galactic dust. Ultraviolet observations of extinction of most SMC stars have shown that dust in the SMC does not absorb in the 2175 Å feature that is so prominent in Milky Way extinction. If aromatic hydrocarbons and featureless extinction curves were ubiquitous in the SMC, then we would conclude that aromatic hydrocarbons are not the carriers of the 2175 Å feature. However, SMC extinction curve measurements are biased toward hot, luminous stars, where aromatic hydrocarbons are destroyed, so that the absence of the 2175 Å bump may not be typical of SMC dust. The presence of aromatic hydrocarbons in the SMC further demonstrates that these molecules exist even in an interstellar medium with an order-of-magnitude lower metallicity than in the disk of the Milky Way.

Accepted by: Astronomy & Astrophysics

For preprints, contact reach@ipac.caltech.edu
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0007382

First Detection of Submillimeter [C i] Emission in the Small Magellanic Cloud

Alberto D. Bolatto^{1}, James M. Jackson^{1}, Kathleen E. Kraemer^{1}, and Xiaolei Zhang^{2}

1 Institute for Astrophysical Research, Boston University, 725 Commonwealth Ave., Boston MA 02215, USA 2 Smithsonian Astrophysical Observatory, 60 Garden St. MS 78, Cambridge, MA 02138, USA

We report the first detection of [C i] (^{3}P_{1} →^{3}P_{0}) emission at 609 μm in a region of the Small Magellanic Cloud (N27). Environments poor in heavy elements and dust such as the SMC are thought to be dominated by photodissociation regions. This is the lowest metallicity source where submillimeter neutral carbon emission has been detected. Studying the [C i]/CO intensity ratio in several sources spanning more than an order of magnitude in metallicity, Z, we find that the [C i]/CO ratio increases for decreasing Z. The existence of such a trend points to a photodissociation origin for most of neutral carbon in molecular clouds, in agreement with standard PDR models. We also report ISO FIR spectroscopic observations of N27, and use them to derive its physical properties. Comparison between the density and radiation field revealed by FIR diagnostics (n ∼ 300—1000 cm^{−3}, χ_{uv} ∼ 30—100 χ_{0}) and those derived from millimeter and submillimeter data (n ∼ 10^{5} cm^{−3}, χ_{uv} < 30 χ_{0}) suggests that the FIR lines originate in more diffuse gas, and are perhaps dominated by the interclump medium. Regardless of the cause, analysis of the FIR and mm-submm data produces a discrepancy of two orders of magnitude for the density of this source.

Accepted by: The Astrophysical Journal Letters For preprints, contact bolatto@bu.edu

Also available from the URLs http://astro.bu.edu/∼bolatto
                          http://arXiv.org/abs/astro-ph/0008055

Carbon in the N159/N160 Complex of the Large Magellanic Cloud

Alberto D. Bolatto^{1}, James M. Jackson^{1}, Frank P. Israel^{2}, Xiaolei Zhang^{3}, & Sungeun Kim^{3,4}

1 Institute for Astrophysical Research, Department of Astronomy, Boston University, 725 Commonwealth Ave., Boston MA 02215, USA 2 Sterrewacht Leiden, P.O. Box 9513, NL 2300-RA Leiden, The Netherlands 3 Smithsonian Astrophysical Observatory, 60 Garden St. MS 78, Cambridge, MA 02138, USA 4 University of Illinois at Urbana-Champaign, Department of Astronomy, 1002 West Green St., Urbana, IL 61801, USA

We present a study of carbon in N159/N160, an H ii region complex in the low metallicity Large Magellanic Cloud. We have mapped this region, which comprises four distinct molecular clouds spanning a wide range of star-formation activity, in four transitions: ^{13}CO (J = 1 → 0), ^{12}CO (J = 2 → 1) and (J = 4 → 3), and [C i] (^{3}P_{1} →^{3}P_{0}). Combining these data with existing [C ii] (^{2}P_{3/2} →^{2}P_{1/2}) observations provides a complete picture of the predominant forms of carbon in the gas phase of the ISM. The new CO (J = 2 → 1) data show that the complex is immersed in an envelope of extended, low level emission, undetected by previous (J = 1 → 0) mapping efforts. The ^{12}CO (J = 2 → 1)/(J = 1 → 0) ratio in this envelope is ∼ 3, a value consistent with optically thin CO emission. The envelope is also relatively bright in [C i] and [C ii], and calculations show that it is mostly photodissociated: it appears to be translucent (A_{v} < 1). Neutral carbon emission in the complex unexpectedly peaks at the quiescent southern cloud (N159S). In the northern portion of the map (the N160 nebula), the H ii regions prominent in [C ii] correspond to holes in the [C i] distribution. Overall we find that, while the I_{[C ii]}/I_{CO} ratio is enhanced with respect to similar complexes in the Milky Way, the I_{[C i]}/I_{CO} ratio appears to be similar or reduced.

Submitted to: The Astrophysical Journal For preprints, contact bolatto@bu.edu

Also available from the URL http://astro.bu.edu/∼bolatto

Compact star clusters of the LMC H ii region N11 C

Mohammad Heydari-Malayeri^{1}, Pierre Royer^{2}, Gregor Rauw^{2} and Nolan R. Walborn^{3}

1 demirm, Observatoire de Paris, 61 Avenue de l’Observatoire, F-75014 Paris, France 2 Institut d’Astrophysique et de Géophysique, Université de Liège, 5, Avenue de Cointe, B-4000 Liège, Belgium 3Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA

Based on imaging and spectroscopy obtained at the ESO NTT telescope and using an efficient image analysis algorithm, we study the core of the LMC OB association LH 13, particularly the two compact stellar clusters Sk–66^{∘}41 and HNT in the H ii region N 11C. We resolve Sk–66^{∘}41 into 15 components and for the first time the HNT cluster into 70 stars, and derive photometry for the members. Moreover, from medium resolution spectroscopy we determine the spectral types for sixteen stars in N 11C. We compare the color-magnitude diagrams of the clusters with that of the field stars and discuss the cluster ages. With an age of ∼ 100 Myr, the HNT cluster appears significantly older than the very young (≤ 5 Myr) Sk–66^{∘}41 starburst. We suggest that most of the ‘field’ O-stars in the core of N 11C have actually been ejected from Sk–66^{∘}41 through dynamical interactions in the compact cluster. The properties of the Sk–66^{∘}41 and HNT clusters suggest that we are viewing different star formation regions lying at different distances along the same line of sight.

Accepted by: Astronomy & Astrophysics

For preprints, contact Mohammad.Heydari-Malayeri@obspm.fr
Also available from astro-ph http://arXiv.org/ps/astro-ph/0007414

WN stars in the LMC: parameters and atmospheric abundances

W.-R. Hamann and L. Koesterke

Astrophysik, Universität Potsdam, Germany

The spectra of 18 WN stars in the Large Magellanic Cloud (LMC) are quantitatively analyzed by means of “standard” Wolf-Rayet model atmospheres, using the helium and nitrogen lines as well as the spectral energy distribution. The hydrogen abundance is also determined. Carbon is included for a subset of 4 stars. The studied sample covers all spectral subtypes (WN2 ... WN9) and also includes one WN/WC transition object.

The luminosities of the program stars span a wide range (log L/L_{⊙} = 5.0 ... 6.5). Due to the given LMC membership, these results are free from uncertainties inferred from the distance. 50% of the studied stars (both, late and early WN subtypes) have rather low luminosity (log L/L_{⊙} ≤ 5.5). This puts tough constraints on their evolutionary formation. If coming from single stars, it provides evidence for strong internal mixing processes.

The empirical mass-loss rates are scaled down by a factor of about two due to the impact of clumping, compared to previous studies adopting homogeneous winds. There is no obvious strong correlation between the mass-loss rates and other parameters like luminosity, temperature and composition.

The stellar parameters for the present LMC sample are not systematically different from those of the Galactic WN stars studied previously with the same techniques, in contrast to the expected metallicity effects.

Accepted by: Astronomy & Astrophysics

For preprints, contact wrh@astro.physik.uni-potsdam.de
Also available from the URL www.astro.physik.uni-potsdam.de

Morphologies and Ages of Star Cluster Pairs and Multiplets in the SMC

M.R. de Oliveira^{1}, C.M. Dutra^{1}, E. Bica^{1} and H. Dottori^{1}

1 Instituto de Fisica-UFRGS, CP 15051, CEP 91501-970 POA - RS, Brazil

An isophotal atlas of 75 star cluster pairs and multiplets in the Small Magellanic Cloud is presented, comprising 176 objects. They are concentrated in the SMC main body. The isophotal contours were made from Digitized Sky Survey images and showed relevant structural features possibly related to interactions in about 25% of the sample. Previous N-body simulations indicate that such shapes could be due to tidal tails, bridges or common envelopes. The diameter ratio between the members of a pair is preferentially in the range 1 − 2, with a peak at 1. The projected separation is in the range ≈ 3 − 22 pc with a pronounced peak at ≈ 13 pc. For 91 objects it was possible to derive ages from Colour-Magnitude Diagrams using the OGLE-II photometric survey. The cluster multiplets in general occur in OB stellar associations and/or HII region complexes. This indicates a common origin and suggests that multiplets coalesce into pairs or single clusters in a short time scale. Pairs in the SMC appear to be mostly coeval and consequently captures are a rare phenomenon. We find evidence that star cluster pairs and multiplets may have had an important role in the dynamical history of clusters presently seen as large single objects.

Accepted by: Astronomy & Astrophysics Supplement For preprints, contact marcio@if.ufrgs.br

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0006369

The distance modulus of the Small Magellanic Cloud based on double-mode Cepheids

G. Kovács^{1}

1Konkoly Observatory, P.O. Box 67, H-1525 Budapest, Hungary

We employ the very recent photometric data of the ogle project together with stellar atmosphere and linear pulsation models to determine the distance modulus of the Small Magellanic Cloud from its double-mode Cepheids. Based on the requirement of obtaining the same distance modulus DM from the two types of variables (fundamental & first overtone and first & second overtone), we get DM = 19.05 mag, with a very small statistical error (standard deviation) of 0.017 mag. Various systematic and zero point ambiguities (primarily those of the color–temperature transformation) lead to an error of ±0.13 mag (estimated 3σ deviation). This result is in very good agreement with the distance modulus of the Large Magellanic Cloud of 18.5 mag, derived earlier from cluster double-mode RR Lyrae stars.

Accepted by: Astronomy & Astrophysics (Letter) For preprints, contact kovacs@konkoly.hu Also available from astro-ph/0007271

Population effects on the Red Giant Clump absolute magnitude, and distance determinations to nearby galaxies

Léo Girardi^{1,3} and Maurizio Salaris^{2,3}

1 Dipartimento di Astronomia, Universitá di Padova, Vicolo dell’Osservatorio 5, I-35122 Padova, Italy 2 Astrophysics Research Institute, Liverpool John Moores University, Twelve Quays House, Egerton Wharf, Birkenhead CH41 1LD, UK 3 Max-Planck-Institut für Astrophysik, Karl-Schwarzschild-Str. 1, D-85741 Garching bei München, Germany

The red giant clump has been recently argued to be a reliable distance indicator for the galaxies in the Local Group. The accuracy of distance determinations based on this method, however, depend on the possible presence of systematic magnitude differences (ΔM_{I}^{RC}) between the local clump revealed by the Hipparcos colour-magnitude diagram, and the clump stars observed in distant galaxies.

In this paper, we re-address the problem of these systematic ‘population’ effects. First, we present tables with the theoretically-predicted I-band clump magnitude as a function of age and metallicity. Simple equations, taken from basic population synthesis theory, are provided for the easy computation of the mean clump magnitude for any given galaxy model. We use our models to explain in some detail what determines the distribution of masses, ages, and metallicities of clump stars in a galaxy. Such an approach has so far been neglected in the analysis of clump data related with distance determinations. We point out that, in galaxies with recent/ongoing star formation (e.g. the disks of spirals), the age distribution of clump stars is strongly biased toward younger (∼ 1 − 3 Gyr) ages, and hence toward higher metallicities. Obviously, this does not happen in galaxies with predominantly old stellar populations (e.g. ellipticals and bulges).

We construct detailed models for the clump population in the local (Hipparcos) sample, the Bulge, Magellanic Clouds, and Carina dSph galaxy. In all cases, star formation rates and chemical enrichment histories are taken from the literature. The Hipparcos model is shown to produce distributions of metallicities, colours, and magnitudes, that are similar to those derived from spectroscopic and Hipparcos data. The Bulge, Magellanic Clouds, and Carina dSph models are used to analyse the values of ΔM_{I}^{RC} for these different stellar systems. We show how the clump–RR Lyrae data from Udalski (1998a) is well reproduced by the models. However, despite the similarity between the models and data, the models indicate that the linear ΔM_{I}^{RC} versus [Fe/H] relations that have been derived from the same data (Udalski 1998a, 2000; Popowski 2000) are not general. In fact, the distribution of clump stars has several factors hidden on it – e.g. the age-metallicity relation, the rate of past star formation – that cannot be described by such relations.

The model behaviour is also supported by empirical data for open clusters by Sarajedini (1999) and Twarog et al. (1999). We argue that Udalski’s (1998b) data for LMC and SMC star clusters do not allow a good assessment of the age dependence of the clump magnitude. Moreover, we remark that similar analyses of cluster data should better include clump stars with ages 1−2 Gyr, which turn out to be very important in determining the mean clump in galaxies with recent star formation.

Finally, we provide revised clump distances to the Bulge, Magellanic Clouds, and Carina dSph, and further comment on their reliability. The largest ΔM_{I}^{RC} values are found for the Magellanic Clouds and Carina dSph, which turn out to be located at distance moduli ∼ 0.2 − 0.3 mag longer than indicated by works which ignore population effects. The Galactic Bulge, instead, is found to be ∼ 0.1 mag closer than indicated by previous works based on the red clump.

Submitted to: Monthly Notices of the Royal Astronomical Society

For preprints, contact lgirardi@pd.astro.it
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0007343

A multi-frequency identification study of the SMC X-ray binary AX J0049-732

M.D. Filipović^{1,2,3}, W. Pietsch^{1} and F. Haberl^{1}

1 Max-Planck-Institut für extraterrestrische Physik, Giessenbachstraße, D-85740 Garching, Germany 2 University of Western Sydney Nepean, P.O. Box 10, Kingswood NSW 2747, Australia 3 Australia Telescope National Facility, CSIRO, P.O. Box 76, Epping NSW 2121, Australia

We present a high-resolution multi-frequency identification study of the Small Magellanic Cloud (SMC) region around the newly discovered X-ray pulsar AX J0049-732. ROSAT PSPC images show three X-ray sources near the ASCA source position. We combine X-ray, radio-continuum and optical data to identify these sources. RX J0049.2-7311 and RX J0049.5-7310 did not show long-term time variability nor pulsations during the ROSAT observations and both are not detected in the radio. We identify RX J0049.5-7310 with an emission line star in the catalogue of Meyssonier & Azzopardi (1993). This catalogue contains several known Be/X-ray binaries strongly suggesting the ROSAT source as new Be/X-ray binary. RX J0049.5-7310 is therefore very likely the counterpart of the ASCA X-ray pulsar AX J0049-732. The third source RX J0049.1-7314 is associated with the supernova remnant (SNR) DEM S49 (B0047-735).

Accepted by: Astronomy & Astrophysics For preprints, contact fica@mpe.mpg.de

Abstracts of Non-Refereed Papers

MACHOs and the clouds of uncertainty

E. Kerins^{1}

1 Theoretical Physics, University of Oxford, 1 Keble Road, Oxford OX1 3NP, UK

I review proposals for explaining the current gravitational microlensing results from the Eros and Macho surveys towards the Magellanic Clouds. Solutions involving massive compact halo objects (MACHOs), both baryonic and non-baryonic, as well as solutions that do not require MACHOs, are discussed. Whilst the existence and nature of MACHOs remains to be established, the prospects for achieving this over the next few years are good.

To appear in: Proceedings of XXth Astrophysics Moriond Meeting: “Cosmological Physics with Gravitational Lensing”, eds. J.-P. Kneib, Y. Mellier, M. Moniez and J. Tran Thanh Van

For preprints, contact e.kerins1@physics.ox.ac.uk
Also available from the URL http://xxx.soton.ac.uk/abs/astro-ph/0007137

Job Opportunity

Postdoctoral Research Associate at the Universidad de Concepción, Concepción, Chile

The Universidad de Concepción has an immediate opening for a Postdoctoral Research Associate to work on a NASA project to establish grid stars for the Space Interferometry Mission (SIM). The position will be for two years, with the possibility of a third year, contingent on NASA funding. The postdoctoral research associate will work with Doug Geisler (Universidad de Concepción and Adjunct Astronomer at the Cerro Tololo Inter-American Observatory) and will reside in Concepción, Chile. The appointment is expected to start 1 October 2000. Starting salary is $30,000. An additional benefit includes an annual trip back to the point-of-origin for the family.

The postdoc will be involved in the preparation, observation, reduction and analysis of photometric and spectroscopic data needed to establish the grid giant stars which will be used by SIM. This database will also be used to study the structure of the Galaxy and to investigate the chemical enrichment history of the halo. A substantial amount of observing at CTIO, ESO and Las Campanas is required. In addition, the postdoc will be encouraged to craft an independent research program in line with his/her research interests. The postdoc will have access to the 10% Chilean share of all the telescopes located in the country, including the VLTs, Gemini South (plus 5% of Gemini North), and Magellan, all progressively coming into operation, in addition to the other telescopes at La Silla, CTIO, and Las Campanas observatories.

Concepción is the second largest city in Chile and lies on the coast 400km south of Santiago, in the lovely south of Chile near many national parks. The Universidad de Concepción is the largest Chilean university outside of Santiago and has recently implemented a strong research program in astronomy

– for more information see http://coma.cfm.udec.cl.

Desirable experience for the successful candidate includes CCD photometric and spectroscopic observational and digital data reduction techniques and general software experience covering UNIX and/or LINUX, IRAF or IDL.

Applicants should send a curriculum vita, a discussion of relevant experience and interest, bibliography and the names and e-mail addresses of three scientists who are familiar with the applicant’s work to the email address listed below. Complete applications should be received by 15 September 2000. AAE/EOE.

Dr. Doug Geisler Universidad de Concepción Grupo de Astronomía Departamento de Física Casilla 160-C Concepción Chile

E-mail: doug@kukita.cfm.udec.cl

Fax: 56-41-224520 Phone: 56-41-203092